# Demystifying amortized causal discovery with transformers

### v2

@@ 4.3 How does CSIvA relate to identifiability theory for causal graphs?
The CSIvA algorithm does not make structural assumptions about the causal model underlying the input data. This implies that the output of this method is unclear: as CSIvA should target the conditional distribution $p(\cdot|\mathcal{D})$ over the space of graphs, in the absence of restrictions on the functional mechanisms and the distribution of the noise terms, the causal graph $X\to Y$ is indistinguishable from $Y\to X$, as they are both equally likely to underlie the joint density $p_{X,Y}$ generating the data. As we discuss in Example 1, the graphical output of the trained architecture could at most identify the equivalence class of the true causal graph. Yet, our experiments of Section 4.2 show that CSIvA is capable of good in-distribution generalization, often inferring the correct DAG at test time. We explain this seeming contradiction with the following hypothesis, which motivates the analysis in the remainder of this section.
Hypothesis 1.
The class of structural causal models that can be identified by CSIvA is defined by the class of structural causal models underlying the generation of the training data.
To support and clarify our statement, we present the following example, adapted from Hoyer et al. (2008).
@@
Consider the causal model $Y=f(X)+N,$ where $f(X)=-X$ and $p_{X},p_{N}$ are Gumbel densities $p_{X}(x)=\operatorname{exp}(-x-\operatorname{exp}(-x))$ and $p_{N}(n)=\operatorname{exp}(-n-\operatorname{exp}(-n))$. This model satisfies the assumptions of the LiNGAM, so it is identifiable, in the sense that a backward linear model with the same distribution does not exist. However, in this special case, we can build a backward nonlinear additive noise model $X=g(Y)+\tilde{N}$ with independent noise terms: taking $p_{Y}(y)=\operatorname{exp}(-y-2\log(1+\operatorname{exp}(-y)))$ to be the density of a logistic distribution, $p_{\tilde{N}}(\tilde{n})=\operatorname{exp}(-2\tilde{n}-\operatorname{exp}(-\tilde{n}))$ and $g(y)=\log(1+\operatorname{exp}(-y))$; we see that $p_{X,Y}$ can factorize according to two opposite causal directions, as $p_{X,Y}(x,y)=p_{N}(y-f(x))p_{X}(x)=p_{\tilde{N}}(x-g(y))p_{Y}(y)$. Given a dataset $\mathcal{D}$ of observations from the forward linear model, causal discovery methods like DirectLiNGAM (Shimizu et al., 2011) can provably identify the correct causal direction $X\rightarrow Y$, assuming that sufficient samples are provided. Instead, the behavior of CSIvA seems hard to predict: given that the network approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the possible graphs, for $\mathcal{D}$ with arbitrary many samples we have $p(X\rightarrow Y|\mathcal{D})=p(Y\rightarrow X|\mathcal{D})=0.5$. On the other hand, given the prior knowledge that the data-generating SCM is a linear non-gaussian additive noise model, we have $p(X\rightarrow Y|\mathcal{D},\textnormal{LiNGAM})=1$, because the LiNGAM is identifiable. In this sense, the class of structural causal models that CSIvA correctly infers appears to be determined by the structural causal models underlying the generation of the training data. Under our Hypothesis 1, training CSIvA exclusively on LiNGAM-generated data is equivalent to learning the distribution $p(\cdot|\mathcal{D},\textnormal{LiNGAM})$, such that the network should be able to identify the forward linear model, whereas it could only infer the equivalence class of the causal graph if its training datasets include observations from a nonlinear additive noise model.
@@
Figure 3: Experiments on identifiability theory. Figure 3(a) shows the SHD of models trained on different ratios of linear and nonlinear invertible data of Example 2. In Figure 3(b) we test the performance on linear-Gaussian data. Models are trained with different ratios of samples from linear and nonlinear SCMs with Gaussian noise terms. The validation results showcase that the networks were trained successfully. In both cases CSIvA behaves according to identifiability theory, failing to predict on invertible data (50:50 ratio) and linear Gaussian models.
@@
Our experiments show that CSIvA learns algorithms that closely follow identifiability theory for causal discovery. In particular, while the method itself does not require explicit assumptions on the data-generating process, the chosen training data ultimately determines the class of causal models identifiable during inference. Notably, previous work has argued that supervised learning approaches in causal discovery would help with "dealing with complex data-generating processes and greatly reduce the need of explicitly crafting identifiability conditions a-priori", Lopez-Paz et al. (2015). In the case of CSIvA, this expectation does not appear to be fulfilled, as the assumptions still need to be encoded explicitly in the training data. However, this observation opens two new important questions: (1) Can we train a single network to encompass multiple (or even all) identifiable causal structures? (2) How much ambiguity might exist between these identifiable models? We start by answering this second question.
@@ 4.4 An identifiability argument in favor of learning from multiple causal models
Example 2 of the previous section shows that elements of distinct classes of identifiable structural causal models, such as LiNGAM and nonlinear ANM, may become non-identifiable when we consider their union. In this section, we discuss the identifiability of the post-additive noise models. Previously, Hoyer et al. (2008) showed that the set of distributions generated according to the additive noise model (3) and that is non-identifiable is negligible. Later, Zhang & Hyvärinen (2009) characterized non-identifiable post-nonlinear models in terms of the properties of their functional mechanisms, and the distribution of the noise terms. In this section, we discuss how these results put together show that the set of distributions generated according to a post-ANM that is non-identifiable is negligible.
@@
$\begin{split}Y=f_{2}(\tilde{Y}),\hskip 10.00002pt\tilde{Y}\coloneqq f_{1}(X)+N_{Y},\\ X=g_{2}(\tilde{X}),\hskip 10.00002pt\tilde{X}\coloneqq g_{1}(Y)+N_{X}.\\ \end{split}$
@@
Equations (7) and (8) reduce the problem of studying the identifiability of a post-ANM to that of studying the identifiability of an additive noise model, as done in Theorem 1 of Hoyer et al. (2008), which we repropose in Appendix E: intuitively, the statement of the theorem says that the space of all continuous distributions generated according to a bivariate additive noise model and that is non-identifiable is contained in a 2-dimensional space. As the space of continuous distributions of random variables is infinite-dimensional, we conclude that the ANM is generally identifiable. Given that, according to Equation (7) and Equation (8), the post-ANM can be refactored in an additive noise model, the guarantees of identifiability still hold (for the formal statement and proof see Appendix E).
@@ 4.5 Can we train CSIvA on multiple causal models for better generalization?
We consider four networks optimized by training of CSIvA on datasets generated from pairs (or triples) of distinct SCMs, with fixed mlp noise and which differ in terms of their mechanisms type: linear and nonlinear; nonlinear and post-nonlinear; linear and post-nonlinear; linear, nonlinear and post-nonlinear. The number of training datasets for each architecture is fixed ($15000$) and equally split between the causal models with different mechanism types. The results of Figure 4 show that the networks trained on mixtures of mechanisms all present significantly better test SHD compared to CSIvA models trained on a single mechanism type. We find that learning on multiple SCMs improves the SHD from $\sim 0.5$ to $\sim 0.2$ both on linear and nonlinear test data (Figures 4(a) and 4(b)), and even better accuracy is achieved on post-nonlinear samples, as shown in Figure 4(c).
@@
Next, we analyze the test performance of three CSIvA networks optimized on samples from structural causal models that have different distributions for their noise terms, while keeping the mechanism types fixed. Figure 5 shows that training over different noises (beta, gamma, gumbel, exponential, mlp, uniform) always results in a network that is agnostic with respect to the noise distributions of the SCM generating the test samples, always achieving SHD $<0.1$, with the exception of datasets with mlp error terms ($0.2$ average SHD on linear and nonlinear data).
@@
We have shown that learning on mixtures of SCMs with different noise term distributions and mechanism types leads to models generalizing to a much broader class of structural causal models during testing. Hence, combining datasets generated from multiple models looks like a promising framework to overcome the limited out-of-distribution generalization abilities of CSIvA observed in Section 4.2. However, it is easier to incorporate prior assumptions on the class of causal mechanisms (linear, non-linear, post-non-linear) compared to the noise distributions (which are potentially infinite). This introduces a trade-off between amortized inference and classical methods for causal discovery: for example, RESIT, NoGAM, and CAM (Peters et al., 2014, Montagna et al., 2023c, Bühlmann et al., 2014) algorithms require no assumptions on the noise type, but only work for a limited class of mechanisms (nonlinear).
@@ 5 Conclusion
In this work, we investigate the interplay between identifiability theory and supervised learning for amortized inference of causal graphs, using CSIvA as the ground of our study. Consistent with classical algorithms, we demonstrate that good performance can be achieved if (i) we have a good prior on the structural causal model generating the test data (ii) the setting is identifiable. In particular, prior knowledge of the test distribution is encoded in the training data in the form of constraints on the structural causal model underlying their generation. With these results, we highlight the need for identifiability theory in modern learning-based approaches to causality, while past works have mostly disregarded this connection. Further, our findings provide the theoretical ground for training on observations sampled from multiple classes of identifiable SCMs, a strategy that improves test generalization to a broad class of causal models. Finally, we highlight an interesting new trade-off regarding identifiability: traditional methods like LiNGAM, RESIT, and PNL require strong restrictions on the structural mechanisms underlying the data generation (linear, nonlinear additive, or post-nonlinear) while generally being agnostic relative to the noise terms distribution. Training on mixtures of causal models instead offers an alternative that is less reliant on assumptions on the mechanisms, while incorporating knowledge about all possible noise distributions in the training data is practically impossible to achieve. We leave it to future work to reproduce our analysis on a wider class of architectures, as well as extending our study to interventional data with more than two nodes.
@@ B.1 Hyperparameters
Num. iterations $<150000$
Batch size $5$
@@ B.2 Synthetic data
We generated datasets from structural causal models with the following distribution of the noise terms: Beta, Gamma, Gaussian (for nonlinear data), Gumbel, Exponential, and Uniform. Additionally, we define the mlp distribution by nonlinear transformations of gaussian samples from a guassian distribution centered at zero and with standard deviation $\sigma$ uniformly sampled between $0.5$ and $1$. The nonlinear transformation is parametrized by a neural network with one hidden layer with $100$ units, and sigmoid activation function. The weights of the network are uniformly sampled in the range $[-1.5,1.5]$. We additionally standardized the output of each mlp sample by the empirical variance computed over all samples.
@@
The nonlinear mechanisms of the PNL model and the nonlinear ANM model are generated by a neural network with one hidden layer with $10$ hidden units, with a parametric ReLU activation function. The network weights are randomly sampled according to a standard Gaussian distribution (we refer to data with nonlinear mechanisms sampled according to this approach as NN-data). The linear mechanisms are generated by sampling the regression coefficients in the range $[-3,-0.5]\cup[0.5,3]$.
@@
We present an extensive list of works adopting neural networks for the sampling of nonlinear mechanisms, similarly to our work: Brouillard et al. (2020; 2021), Lippe et al. (2022), Bello et al. (2022), Montagna et al. (2023a; b), Ke et al. (2023a; b), Reizinger et al. (2023), Massidda et al. (2023), Tran et al. (2024). This suggests that our data generation strategy is established in the literature of causality. Additional experiments with sampling of nonlinear mechanisms from Gaussian processes are presented in Appendix D.7.
@@ Appendix C CSIvA identifiability properties on multivariate SCMs
The above theorem states that a restricted ANM is identifiable. According to our Definition 2, an additive noise model $X_{h}\coloneqq f_{h}(X_{\operatorname{PA}^{\mathcal{G}}_{h}})+N_{h},h=1,\ldots,k$ is restricted if each pair of connected nodes $X_{i}\rightarrow X_{j}$, for each $X_{\operatorname{PA}^{\mathcal{G}}_{j}}\setminus\{X_{i}\}\subseteq X_{S}\subseteq X_{\operatorname{ND}_{j}}^{\mathcal{G}}\setminus\{X_{i},X_{j}\}$ (think of $X_{S}$ as the set of all possible causes of $X_{j}$, except $X_{i}$), can define a bivariate SCM of the form $(f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},\cdot),p_{X_{i}|X_{S}=x_{S}},p_{N_{j}})$ that satisfies Condition 1, i.e. that is identifiable. How does this relate to our findings? Our experiments and analysis of Section 4.3 validate the hypothesis that transformers align with the theory of identifiability in the case of training and inference on bivariate graphs. Given Theorem 1, we know that multivariate identifiability is a property of SCMs where each pair of causes and effects can define a bivariate structural causal model that is itself identifiable: this implies that the empirical guarantees of identifiability we verify for transformers (via CSIvA) on bivariate models must extend to multivariate models. This is apparent by contradiction: say we train a CSIvA architecture that can infer the causal direction of a multivariate linear Gaussian model (which is notoriously non-identifiable). This means that our algorithm can infer the causal direction for each bivariate subgraph consisting of two variables connected according to a linear Gaussian structural equation: this would contradict our experimental results presented in Fig. 3(b) and analyzed in Section 4.3.

### v3

@@ 4.3 How does CSIvA relate to identifiability theory for causal graphs?
The CSIvA algorithm does not make structural assumptions about the causal model underlying the input data. This implies that the output of this method is unclear: as CSIvA should target the conditional distribution $p(\cdot|\mathcal{D})$ over the space of graphs, in the absence of restrictions on the functional mechanisms and the distribution of the noise terms, the causal graph $X\to Y$ is indistinguishable from $Y\to X$, as they are both equally likely to underlie the joint density $p_{X,Y}$ generating the data. As we discuss in Example 1, the graphical output of the trained architecture could at most identify the equivalence class of the true causal graph. Yet, our experiments of Section 4.2 show that CSIvA is capable of good in-distribution generalization, often inferring the correct DAG at test time. We explain this seeming contradiction with the following hypothesis, which motivates the experimental analysis in the remainder of this section.
Hypothesis (informal). At test time, the datasets on which CSIvA is expected to generalize well are determined by the family of SCMs on which training occurred, and can be analyzed using the tools from identifiability theory.
Notably, if this hypothesis is verified, we can analyse when CSIvA is expected or not to work well; the remainder of this work empirically studies this claim. Before moving forward, we present the following example adapted from Hoyer et al. (2008), which supports and clarifies our statement.
@@
Consider the causal model $Y=f(X)+N,$ where f(X)=Xf(X)=-X and $p_{X},p_{N}$ are Gumbel densities $p_{X}(x)=\operatorname{exp}(-x-\operatorname{exp}(-x))$ and $p_{N}(n)=\operatorname{exp}(-n-\operatorname{exp}(-n))$. This model satisfies the assumptions of the LiNGAM, so it is identifiable, in the sense that a backward linear model (i.e., $Y\to X$ with linear mechanism $f$) with the same distribution does not exist. However, in this special case, we can build a backward nonlinear additive noise model $X=g(Y)+\tilde{N}$ with independent noise terms: taking $p_{Y}(y)=\operatorname{exp}(-y-2\log(1+\operatorname{exp}(-y)))$ to be the density of a logistic distribution, $)p_{\tilde{N}}(\tilde{n})=\operatorname{exp}(-2\tilde{n}-\operatorname{exp}(-\tilde{n}$)) and $g(y)=\log(1+\operatorname{exp}(-y))$; we see that $p_{X,Y}$ can factorize according to two opposite causal directions, as $p_{X,Y}(x,y)=p_{N}(y-f(x))p_{X}(x)=p_{\tilde{N}}(x-g(y))p_{Y}(y)$. Given a dataset $\mathcal{D}$ of observations from the forward linear model $X\to Y$, causal discovery methods like DirectLiNGAM (Shimizu et al., 2011) can provably identify the correct causal direction ($X\rightarrow Y$), assuming that sufficient samples are provided. Instead, the behavior of CSIvA seems hard to predict: given that the network approximates the conditional distribution $p(\cdot|\mathcal{D})$ over the possible graphs, for $\mathcal{D}$ with arbitrary many samples we have $p(X\rightarrow Y|\mathcal{D})=p(Y\rightarrow X|\mathcal{D})=0.5$. On the other hand, given the prior knowledge that the data-generating SCM is a linear non-gaussian additive noise model, we have $p(X\rightarrow Y|\mathcal{D},\textnormal{LiNGAM})=1$, because the LiNGAM is identifiable. In this sense, the class of structural causal models that CSIvA correctly infers appears to be determined by the structural causal models underlying the generation of the training data. Under this hypothesis, training CSIvA exclusively on LiNGAM-generated data is equivalent to learning the distribution $p(\cdot|\mathcal{D},\textnormal{LiNGAM})$, such that the network should be able to identify the forward linear model, whereas it could only infer the equivalence class of the causal graph if its training datasets include observations from a nonlinear additive noise model.
@@
Figure 3: Experiments on identifiability theory. Figure 3(a) shows the SHD of models trained on different ratios of linear and nonlinear invertible data of Example 2. In Figure 3(b) we test the performance on linear-Gaussian data. Models are trained with different ratios of samples from linear and nonlinear SCMs with Gaussian noise terms. The validation results showcase that the networks were trained successfully. In both cases, CSIvA behaves according to identifiability theory, failing to predict on invertible data (50:50 ratio) and linear Gaussian models.
@@
Our experiments show that CSIvA learns algorithms that closely follow identifiability theory for causal discovery. In particular, while the method itself does not require explicit assumptions on the data-generating process, the chosen training data ultimately determines the class of causal models identifiable during inference. Notably, previous work has argued that supervised learning approaches in causal discovery would help with "dealing with complex data-generating processes and greatly reduce the need of explicitly crafting identifiability conditions a-priori", Lopez-Paz et al. (2015). In the case of CSIvA, this expectation does not appear to be fulfilled, as the assumptions still need to be encoded explicitly in the training data. However, this observation opens two new important questions: (1) Can we train a single network to correctly classify multiple (or even all) identifiable causal structures? (2) How much ambiguity might exist between these identifiable models? We start by answering this second question.
@@ 4.4 An identifiability argument in favor of learning from multiple causal models
Example 2 of the previous section shows that elements of distinct classes of identifiable structural causal models, such as LiNGAM and nonlinear ANM, may become non-identifiable when we consider their union. In this section, we discuss the identifiability of the post-additive noise models. Previously, Hoyer et al. (2008) showed that the set of distributions generated according to the additive noise model equation 3 and that is non-identifiable is negligible. Later, Zhang and Hyvärinen (2009) characterized non-identifiable post-nonlinear models in terms of the properties of their functional mechanisms, and the distribution of the noise terms. In this section, we discuss how these results put together show that the set of distributions generated according to a post-ANM that is non-identifiable is negligible.
@@
[…]
@@
Equations (7) and (8) reduce the problem of studying the identifiability of a post-ANM to that of studying the identifiability of an additive noise model, as done in Theorem 1 of Hoyer et al. (2008), which we repropose in Appendix˜E: intuitively, the statement of the theorem says that the space of all continuous distributions generated according to a bivariate additive noise model and that is non-identifiable is contained in a 2-dimensional space. As the space of continuous distributions of random variables is infinite-dimensional, we conclude that the ANM is generally identifiable. Given that, according to Equation equation 7 and Equation equation 8, the post-ANM can be refactored in an additive noise model, the guarantees of identifiability still hold (for the formal statement and proof see Appendix˜E).
@@ 4.5 Can we train CSIvA on multiple causal models for better generalization?
We consider four networks optimized by training of CSIvA on datasets generated from pairs (or triples) of distinct SCMs, with fixed mlp noise and which differ in terms of their mechanisms type: linear and nonlinear; nonlinear and post-nonlinear; linear and post-nonlinear; linear, nonlinear and post-nonlinear. The number of training datasets for each architecture is fixed ($15000$) and equally split between the causal models with different mechanism types. The results of Figure 4 show that the networks trained on mixtures of mechanisms all present significantly better test SHD compared to CSIvA models trained on a single mechanism type. We find that learning on multiple SCMs improves the SHD from $\sim\hskip-1.42262pt0.5$ to $\sim\hskip-1.42262pt0.2$ both on linear and nonlinear test data (Figures 4(a) and 4(b)), and even better accuracy is achieved on post-nonlinear samples, as shown in Figure 4(c).
@@
Next, we analyze the test performance of three CSIvA networks optimized on samples from structural causal models that have different distributions for their noise terms, while keeping the mechanism types fixed. Figure 5 shows that training over different noises (beta, gamma, gumbel, exponential, mlp, uniform) always results in a network that is agnostic with respect to the noise distributions of the SCM generating the test samples, always achieving SHD <0.1<0.1, with the exception of datasets with mlp error terms ($0.2$ average SHD on linear and nonlinear data).
@@
We have shown that learning on mixtures of SCMs with different noise term distributions and mechanism types leads to models generalizing to a much broader class of structural causal models during testing. Hence, combining datasets generated from multiple models looks like a promising framework to overcome the limited out-of-distribution generalization abilities of CSIvA observed in Section 4.2. However, it is easier to incorporate prior assumptions on the class of causal mechanisms (linear, non-linear, post-non-linear) compared to the noise distributions (which are potentially infinite). This introduces a trade-off between amortized inference and classical methods for causal discovery: for example, RESIT, NoGAM, AdaScore, and LNMIX (Peters et al., 2014; Montagna et al., 2023c; 2025; Liu et al., 2024) algorithms require no assumptions on the noise type, but only work for a limited class of mechanisms (linear and nonlinear with additive noise).
@@ 5 Conclusion
In this work, we investigate the interplay between identifiability theory and supervised learning for amortized inference of causal graphs, using CSIvA as the ground of our study. Consistent with classical algorithms, we demonstrate that good performance can be achieved if (i) we have a good prior on the structural causal model generating the test data (ii) the setting is identifiable. In particular, prior knowledge of the test distribution is encoded in the training data in the form of constraints on the structural causal model underlying their generation. With these results, we highlight the need for identifiability theory in modern learning-based approaches to causality, while past works have mostly disregarded this connection. Further, our findings provide the theoretical ground for training on observations sampled from multiple classes of identifiable SCMs, a strategy that improves test generalization to a broad class of causal models. Finally, we highlight an interesting new trade-off regarding identifiability: traditional methods like LiNGAM, RESIT, and PNL require strong restrictions on the structural mechanisms underlying the data generation (linear, nonlinear additive, or post-nonlinear) while generally being agnostic relative to the noise terms distribution. Training on mixtures of causal models instead offers an alternative that is less reliant on assumptions on the mechanisms, while incorporating knowledge about all possible noise distributions in the training data is practically impossible to achieve. We leave it to future work to reproduce our analysis on a wider class of architectures, as well as extend our study to interventional data with more than two nodes.
@@ B.1 Hyperparameters
Num. iterations <150000<150000
Batch size 55
@@ B.2 Synthetic data
We generated datasets from structural causal models with the following distribution of the noise terms: Beta, Gamma, Gaussian (for nonlinear data), Gumbel, Exponential, and Uniform. Additionally, we define the mlp distribution by nonlinear transformations of gaussian samples from a guassian distribution centered at zero and with standard deviation $\sigma$ uniformly sampled between $0.5$ and $1$. The nonlinear transformation is parametrized by a neural network with one hidden layer with $100$ units, and sigmoid activation function. The weights of the network are uniformly sampled in the range [1.5,1.5][-1.5,1.5]. We additionally standardized the output of each mlp sample by the empirical variance computed over all samples.
@@
The nonlinear mechanisms of the PNL model and the nonlinear ANM model are generated by a neural network with one hidden layer with $10$ hidden units, with a parametric ReLU activation function. The network weights are randomly sampled according to a standard Gaussian distribution (we refer to data with nonlinear mechanisms sampled according to this approach as NN-data). The linear mechanisms are generated by sampling the regression coefficients in the range $3,-0.5]\cup[0.5,3]$.
@@
We present an extensive list of works adopting neural networks for the sampling of nonlinear mechanisms, similarly to our work: Brouillard et al. (2020; 2021); Lippe et al. (2022); Bello et al. (2022); Montagna et al. (2023a; b); Ke et al. (2023a; b); Reizinger et al. (2023); Massidda et al. (2023); Tran et al. (2024). This suggests that our data generation strategy is established in the literature of causality. Additional experiments with sampling of nonlinear mechanisms from Gaussian processes are presented in Appendix D.7.
@@ Appendix C CSIvA identifiability properties on multivariate SCMs
The above theorem states that a restricted ANM is identifiable. According to our Definition 2, an additive noise model $X_{h}\coloneqq f_{h}(X_{\operatorname{PA}^{\mathcal{G}}_{h}})+N_{h},h=1,\ldots,k$ is restricted if each pair of connected nodes $X_{i}\rightarrow X_{j}$, for each $X_{\operatorname{PA}^{\mathcal{G}}_{j}}\setminus\{X_{i}\}\subseteq X_{S}\subseteq X_{\operatorname{ND}_{j}}^{\mathcal{G}}\setminus\{X_{i},X_{j}\}$ (think of $X_{S}$ as the set of all possible causes of $X_{j}$, except $X_{i}$), can define a bivariate SCM of the form $(f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},\cdot),p_{X_{i}|X_{S}=x_{S}},p_{N_{j}})$ that satisfies Condition 1, i.e. that is identifiable. How does this relate to our findings? Our experiments and analysis of Section 4.3 validate the hypothesis that transformers align with the theory of identifiability in the case of training and inference on bivariate graphs. Given Theorem 1, we know that multivariate identifiability is a property of SCMs where each pair of causes and effects can define a bivariate structural causal model that is itself identifiable: this implies that the empirical guarantees of identifiability we verify for transformers (via CSIvA) on bivariate models must extend to multivariate models. This is apparent by contradiction: say we train a CSIvA architecture that can infer the causal direction of a multivariate linear Gaussian model (which is notoriously non-identifiable). This means that our algorithm can infer the causal direction for each bivariate subgraph consisting of two variables connected according to a linear Gaussian structural equation: this would contradict our experimental results presented in Fig.˜3(b) and analyzed in Section˜4.3.